OpenStudy (anonymous):

The profit function p(x) of a tour operator is modeled by p(x) = −2x2 + 700x − 10000 where x is the average number of tours he arranges per day. What is the range of the average number of tours he must arrange per day to earn a monthly profit of at least $50,000? between 150 and 200, exclusive more than 200 between 150 and 200, inclusive less than 150

OpenStudy (mathmate):

hint: p(x) means replace "x" on the right hand side by the value of x. For example, \(p(100)=-2(100^2)+700(100)-10000=-20000+70000-10000=40000\)

OpenStudy (mathmate):

By the way, the question should have read The \(monthly\) profit function p(x)..... to avoid confusion with the "number of tours he arranges per \(day\).

OpenStudy (anonymous):

I got 40,000 from the equation you gave me

OpenStudy (anonymous):

what do I have to do next?.,.

OpenStudy (mathmate):

The number 40000 equals f(100), which is an example of how to calculate f(x) for x=100. You need to use other values of x to solve the problem. Read the question again to find more hints. :)